# A Comprehensive Study of Shapley Value in Data Analytics [Experiment, Analysis & Benchmark]

### v6

@@ Abstract.
Over the recent years, Shapley value (SV), a solution concept from cooperative game theory, has found numerous applications in data analytics (DA). This paper provides the first comprehensive study of SV used throughout the DA workflow, clarifying the key variables in defining DA-applicable SV and the essential functionalities that SV can provide for data scientists. We condense four primary challenges of using SV in DA, namely computation efficiency, approximation error, privacy preservation, and interpretability, then disentangle the resolution techniques from existing arts in this field, analyze and discuss the techniques w.r.t. each challenge and the potential conflicts between challenges. We also implement SVBench, a modular and extensible open-sourced framework for developing SV applications in different DA tasks, and conduct extensive evaluations to validate our analyses and discussions. Based on the qualitative and quantitative results, we identify the limitations of current efforts for applying SV to DA and highlight the directions of future research and engineering.
@@
PVLDB, 14(1): 1-12, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Recently, data scientists have applied Shapley value (SV), a method derived from cooperative game theory for fairly distributing the total gains generated by the coalition of all players (Shapley, 1953), to numerous tasks throughout the DA workflow. Figure 1 depicts a series of examples applying SV to the tasks analyzing medical images, where the analytical objects include pixels, images, image sets, ML models, etc. The application purposes can be summarized into four categories: (1) pricing, to determine the net worth of analytical objects for trading, such as buying image datasets in the DC task and selling well-trained models in the RT task; (2) selection, to select qualified and importance analytical objects for exploration, for example, selecting pixels important to reduce training losses to learn models in the DO task; (3) weighting, to assign reasonable weights to analytical objects collected from multiple sources for valid fusion of those objects, for instance, weighting local models collected in the DM task, where federated learning (FL) (McMahan et al., 2016; Liu et al., 2022a) is used to protect privacy within medical images, for fusing as a valid global model whose accuracy exceeds a certain threshold; (4) attribution, to explain data exploration outputs, e.g., explaining how pixels impact the model predictions in the RI task.
@@
Table 1. Comparison with related works. Our work focuses on the techniques (or called building blocks of SV applications/algorithms) solving four challenges of SV: computation efficiency (eff.), approximation error (err.), privacy preservation (priv.), and interpretability (int.).
To better understand the usage of SV in the DA domain, several surveys have been proposed (Rozemberczki et al., 2022; Chen et al., 2023; Li et al., 2024; Bertossi et al., 2023; Sim et al., 2022; Pei, 2020; Tang et al., 2020; Olsen et al., 2024). However, they do not cover the entire life cycle of DA, leaving a gap in holistic guidance for applying SV to diverse DA tasks. As summarized in Table 1, prior works predominantly concentrated on ad-hoc SV implementations, mainly addressing computation efficiency instead of other challenges such as interpretability. Furthermore, they analyze SV applications/computing algorithms as monolithic units rather than decomposing them into reusable building blocks, which is a limitation that stifles the development of modular, extensible frameworks from a DA application perspective. For instance, existing tools like SHAP (Lundberg and Lee, 2017a), tailored to RI tasks, and DataShapley (Ghorbani and Zou, 2019), bound to DV tasks, are constrained by rigid task-specific configurations, sacrificing flexibility and universality. These gaps leave practitioners ill-equipped to navigate SV’s foundational design principles, such as cooperative game modeling, or resolve conflicting challenges (e.g., approximation error vs. efficiency) when adapting SV to new DA tasks.
@@
(1) We present the first comprehensive survey of SV applied throughout the DA workflow, clarify the key variables in defining DA-applicable SV, and reveal the essential functionalities that SV can provide for DA. We also condense four technical challenges of using SV in DA, analyze and discuss the building blocks of solutions w.r.t. each challenge and potential conflicts between challenges. (§3)
(2) We propose SVBench, a modular and extensible open-sourced framework for developing SV applications. SVBench integrates abundant techniques addressing key challenges of using SV in DA and allows a flexible usage of these techniques to build up new algorithms. The developers, as well as the academicians, could extend the framework with flexible APIs for succeeding studies on SV. (§4)
(3) We conduct extensive evaluations to consolidate our analyses and discussions, involving possible technical combinations for efficient computation, trade-offs in handling different challenges, and key factors affecting mainstream SV interpretations. Experiment results also validate the usability and modularity of SVBench. Through experiments, we reveal limiting factors of existing efforts and highlight future research and engineering directions. (§5)
@@ 2. Preliminaries
Eq. 1 and Eq. 2 show the mathematical formulation of SV from different perspectives. Eq. 1 is given based on Definition 2, while Eq. 2 is given from the perspective of expectation calculation. The two formulas are interchangeable in SV computation. However, Eq. 2 would be preferred if a developer expects to compute SV using expectation calculation techniques in mathematical statistics (even if not originally devised for SV, e.g., Monte Carlo methods). In Eq. 2, $\pi(\mathcal{N})$ is the set of all possible permutations of players in the grand coalition, and $P(\mathcal{O},p_{i})$ is the set of predecessors of player $p_{i}$ in a specific permutation $\mathcal{O}\in\pi(\mathcal{N})$.
@@ 3. The Shapley Value in Data Analytics
Cooperative Game Modeling Player Data Feature Data Tuple Dataset Data Derivative Utility Goodness -of-fit Task Output Goodness -of-fit Task Output Goodness -of-fit Task Output Goodness -of-fit Task Output Purpose Dataset PricingFeature SelectionFeature AttributionFeature SelectionDataset PricingTuple SelectionFeature AttributionTuple PricingTuple SelectionTuple AttributionTuple AttributionDataset PricingDataset SelectionDataset WeightingDataset AttributionDataset AttributionModel PricingModel SelectionModel WeightingModel Weighting Task & Reference DC(Azcoitia et al., 2023)DO(Chu and Chan, 2020; Fryer et al., 2021; Cohen et al., 2005; Meepaganithage et al., 2024; Dong et al., 2021; Yin et al., 2022; Tripathi et al., 2020; Cohen et al., 2007; Pelegrina et al., 2024a; Zöllner and Gürtler, 2023)RI(Watson et al., 2023; Pelegrina and Siraj, 2024; Pelegrina et al., 2024b; Kolpaczki et al., 2024b; Beechey et al., 2023; Shi et al., 2024a, b)DO(Sebastián and González-Guillén, 2024; Marcílio and Eler, 2020; Kim et al., 2024)DM(Wang et al., 2019)DM(Kara et al., 2022; Huber et al., 2022)RI(Wang et al., 2024a; Olsen et al., 2022; Merrick and Taly, 2020; Redelmeier et al., 2020; Aas et al., 2021b, a; Frye et al., 2021; Kumar et al., 2021b; Covert and Lee, 2020; Mitchell et al., 2022; Jethani et al., 2021; Wang et al., 2023; Luo et al., 2022)(Kumar et al., 2020; Duval and Malliaros, 2021; Kwon and Zou, 2022; Zheng et al., 2022b; Demertzis et al., 2022; Cuzzocrea et al., 2023; Sundararajan and Najmi, 2020; Kolpaczki et al., 2024a; Bento et al., 2021; Ma and Tourani, 2020; Williamson and Feng, 2020; Olsen et al., 2024; Huang and Marques-Silva, 2024)DC(Azcoitia et al., 2022; Tian et al., 2022b; Zhang et al., 2023b; Wu et al., 2023b; Xia et al., 2025)DV(Jia et al., 2019; Zhang et al., 2023a; Shobeiri and Aajami, 2021; Xu et al., 2023; Ghorbani et al., 2020; Kwon et al., 2021; Sun et al., 2024)DV(Schoch et al., 2022; Shobeiri and Aajami, 2022; Xia et al., 2024; Ghorbani and Zou, 2019; Wang et al., 2025)DM(Courtnage, 2022; Ghorbani et al., 2022; Shim et al., 2021; Courtnage and Smirnov, 2021; Pandl et al., 2021)RI(Heuillet et al., 2022; Ruggeri et al., 2024; Li et al., 2021)RI(Bertossi et al., 2023; Deutch et al., 2022; Livshits et al., 2021b, a)DC(Luo et al., 2024; Xia et al., 2023; Liu et al., 2021; Agarwal et al., 2019; Tian et al., 2022a; Si and Pei, 2024; Sim et al., 2020; Nguyen et al., 2022; Sim et al., 2020)DV(Liu et al., 2023; Sim et al., 2022; Wang et al., 2024b; Garrido et al., [n.d.])DM(Zhang et al., 2020)RI(Farchi et al., 2021)RI(Brandsæter and Glad, 2024)RT(Xu et al., 2024)DM(Zheng et al., 2022a; Yang et al., 2023b, 2022a; Fan et al., 2022; Yang et al., 2022b, 2024; Batool et al., 2022; Liu et al., 2022a; Arbaoui et al., 2024; Liu et al., 2020; Tastan et al., 2024; Xu et al., 2021)(Lin et al., 2023; Yu et al., 2020a; Ma et al., 2021; Huang et al., 2020; Shi et al., 2022; Yu et al., 2020b; Wei et al., 2020; Liu et al., 2022b; Wang et al., 2022a; Zhu et al., 2023b; You et al., 2021; Cheng et al., 2024; Wang et al., 2022b, 2024c)DM(Guan et al., 2022; Wang et al., 2022b; Huang et al., 2021; Otmani et al., 2024; Xi et al., 2021; Singhal et al., 2024; Wang et al., 2020; Nagalapatti and Narayanam, 2021; Salarbashishahri et al., 2022; Rozemberczki and Sarkar, 2021)DM(Wu et al., 2023a; Salarbashishahri et al., 2022; Sun et al., 2023; Tang et al., 2021; Yang et al., 2023a; Franses et al., 2024; Drungilas et al., 2023)DM(Bimonte et al., 2024)❶❷❶❷❶❷❶❷❶❷❶❷❶❷
@@ 3.1. Cooperative Game Modeling
Finding 1. Players may differ from the analytical object. While players in SV typically align with the analytical object of a DA task, they can also represent sub-components of the analytical object for specialized objectives.
@@
In contrast, Wang et al. (2019) defined players as the sub-components of the analytical object in a multi-group data valuation task. The task is to evaluate contributions to Cervical cancer prediction across five disjoint groups, where each group’s dataset contains three unique data features (e.g., age) with no feature overlap between groups. Here, the player is the data feature, a sub-component of the core analytical object (each group’s dataset). The flexibility in defining players as either the analytical object or its sub-components underscores the adaptability to DA task-specific interpretability needs.
@@
The utility function has two kinds of outputs: (1) the goodness-of-fit score between the DA task’s outputs and ground truth facts, including but not limited to test accuracy, training loss, confidence scores, etc., (2) the task outputs themselves, more specifically, predictive outputs from the ML models learned in DA tasks or the answers to queries/questions on the data used in the DA workflow.
@@
Finding 2. Utility definitions are task-dependent. The choice of utility is dictated by the objectives of the DA task: when maximizing/minimizing goodness-of-fit scores (e.g., accuracy), the utility directly reflects those scores; conversely, when explaining task outputs or quantifying data impacts, the utility is defined by the outputs themselves.
@@
Discussion. Figure 1 exemplifies this duality. In the DV task, utility is tied to test accuracy to optimize a medical image classifier, while in the RI task, utility derives from predictive outputs to explain how pixel-level variations influence results. Notably, a single player-utility pairing, such as defining players as ML models and utility as test accuracy, can serve multiple DA tasks (e.g., DM and RT in the figure). Similarly, one task may leverage SV for diverse purposes: the DM task, for instance, can use SV both to weight local models and assign pricing incentives, ensuring higher model quality while aligning contributions with economic rewards. This adaptability reveals SV’s capacity to address heterogeneous analytical goals through context-aware utility design.
@@
Finding 3. The real-world functionalities of applying SV lie in three aspects: (1) to construct fair marketplaces for data and its related products, incentivizing data sharing; (2) to improve the quality of DA outputs while reducing economic cost; (3) to transfer DA outputs into actions to solve real-world problems.
@@
Discussion. Figure 1 exemplifies the three functionalities that different DA tasks target to achieve. DC and RT tasks seek to provide helpful results to construct fair marketplaces for data in the medical domain. DV and DM tasks strive to improve the test accuracy of a medical image classification model while reducing learning cost. The RI task intends to transfer predictions from a medical image classification model into actions assisting diagnosis. As summarized in Figure 2, to serve the first functionality, the purpose of applying SV is typically pricing. For the second functionality, the purposes of applying SV include pricing, selection, and weighting. To fulfill the last functionality, the purpose of applying SV is primarily attribution. Demanders can design SV applications that satisfy their specific requirements according to Figure 2 and Findings 1-3.

### v7

@@ Abstract.
Over the recent years, Shapley value (SV), a solution concept from cooperative game theory, has found numerous applications in data analytics (DA). This paper presents the first comprehensive study of SV used throughout the DA workflow, clarifying the key variables in defining DA-applicable SV and the essential functionalities that SV can provide for data scientists. We condense four primary challenges of using SV in DA, namely computation efficiency, approximation error, privacy preservation, and interpretability, disentangle the resolution techniques from existing arts in this field, then analyze and discuss the techniques w.r.t. each challenge and the potential conflicts between challenges. We also implement SVBench, a modular and extensible open-source framework for developing SV applications in different DA tasks, and conduct extensive evaluations to validate our analyses and discussions. Based on the qualitative and quantitative results, we identify the limitations of current efforts for applying SV to DA and highlight the directions of future research and engineering.
@@
PVLDB, 18(9): 3077 - 3092, 2025.
doi:10.14778/3746405.3746429 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 9 ISSN 2150-8097.
doi:10.14778/3746405.3746429
@@ 1. Introduction
Recently, data scientists have applied Shapley value (SV), a method derived from cooperative game theory for fairly distributing the total gains generated by the coalition of all players (Shapley, 1953), to numerous tasks throughout the DA workflow. Figure 1 depicts a series of examples applying SV to the tasks analyzing medical images, where the analytical objects include pixels, images, image sets, ML models, etc. The application purposes can be summarized into four categories: (1) pricing, to determine the net worth of analytical objects for trading, such as buying image datasets in the DC task and selling well-trained models in the RT task; (2) selection, to select qualified and important analytical objects for exploration, for example, selecting pixels important to reduce training losses to learn models in the DO task; (3) weighting, to assign reasonable weights to analytical objects collected from multiple sources for valid fusion of those objects, for instance, weighting local models collected in the DM task, where federated learning (FL) (McMahan et al., 2016; Liu et al., 2022a) is used to protect privacy within medical images, for fusing as a valid global model whose accuracy exceeds a certain threshold; (4) attribution, to explain data exploration outputs, e.g., explaining how pixels impact the model predictions in the RI task.
To better understand the usage of SV in the DA domain, several surveys have been proposed (Rozemberczki et al., 2022; Chen et al., 2023; Li et al., 2024; Bertossi et al., 2023; Sim et al., 2022; Pei, 2020; Tang et al., 2020; Olsen et al., 2024). However, they do not cover the entire life cycle of DA, leaving a gap in a holistic guidance for applying SV to diverse DA tasks. As summarized in Table 1, prior works predominantly concentrated on ad-hoc SV implementations, mainly addressing computation efficiency instead of other challenges such as interpretability. Furthermore, they analyze SV applications or computing algorithms as monolithic units rather than decomposing them into reusable building blocks, which is a limitation that stifles the development of modular extensible frameworks from a DA application perspective. For instance, existing tools like SHAP (Lundberg and Lee, 2017a) (tailored to RI tasks) and DataShapley (Ghorbani and Zou, 2019) (bound to DV tasks) are constrained by rigid task-specific configurations, sacrificing flexibility and universality. These gaps leave practitioners ill-equipped to navigate SV’s foundational design principles (such as cooperative game modeling) or resolve conflicting challenges (e.g., approximation error vs. efficiency) when adapting SV to new DA tasks.
@@
Table 1. Comparison with related works. Our work focuses on the techniques (or called building blocks of SV applications and algorithms) solving four challenges of SV: computation efficiency (eff.), approximation error (err.), privacy preservation (priv.), and interpretability (int.).
@@
  • We present the first comprehensive survey of SV applied throughout the DA workflow, clarify the key variables in defining DA-applicable SV, and reveal the essential functionalities that SV can provide for DA. We also condense four technical challenges of using SV in DA, analyze and discuss the building blocks of solutions w.r.t. each challenge and potential conflicts between challenges. (§3)
  • We propose SVBench, a modular and extensible open-source framework for developing SV applications. SVBench integrates abundant techniques addressing key challenges of using SV in DA and allows a flexible usage of these techniques to build up new algorithms. The developers, as well as the academicians, could extend the framework with flexible APIs for succeeding studies on SV. (§4)
  • We conduct extensive evaluations to consolidate our analyses and discussions, involving possible technical combinations for efficient computation, trade-offs in handling different challenges, and key factors affecting mainstream SV interpretations. Experiment results also validate the usability and modularity of SVBench. Through experiments, we reveal limiting factors of existing efforts and highlight future research and engineering directions. (§5)
@@ 2. Preliminaries
Equations 1 and 2 show the mathematical formulation of SV from different perspectives. Equation 1 is given based on Definition 2, while Equation 2 is given from the perspective of expectation calculation. The two formulas are interchangeable in SV computation. However, Equation 2 would be preferred if a developer expects to compute SV using expectation calculation techniques in mathematical statistics (even if not originally devised for SV, e.g., Monte Carlo methods). In Equation 2, $\pi(\mathcal{N})$ is the set of all possible permutations of players in the grand coalition, and $P(\mathcal{O},p_{i})$ is the set of predecessors of player $p_{i}$ in a specific permutation $\mathcal{O}\in\pi(\mathcal{N})$.
@@ 3. The Shapley Value in Data Analytics
Cooperative Game Modeling Player Data Feature Data Tuple Dataset Data Derivative Utility Goodness -of-fit Task Output Goodness -of-fit Task Output Goodness -of-fit Task Output Goodness -of-fit Task Output Purpose Dataset PricingFeature SelectionFeature AttributionFeature SelectionDataset PricingTuple SelectionFeature AttributionTuple PricingTuple SelectionTuple AttributionTuple AttributionDataset PricingDataset SelectionDataset WeightingDataset AttributionDataset AttributionModel PricingModel SelectionModel WeightingModel Weighting Task & Reference DC(Azcoitia et al., 2023)DO(Chu and Chan, 2020; Fryer et al., 2021; Cohen et al., 2005; Meepaganithage et al., 2024; Dong et al., 2021; Yin et al., 2022; Tripathi et al., 2020; Cohen et al., 2007; Pelegrina et al., 2024a; Bosker et al., 2024)RI(Watson et al., 2023; Pelegrina and Siraj, 2024; Pelegrina et al., 2024b; Kolpaczki et al., 2024b; Beechey et al., 2023; Shi et al., 2024a, b)DO(Sebastián and González-Guillén, 2024; Marcílio and Eler, 2020; Kim et al., 2024)DM(Wang et al., 2019)DM(Kara et al., 2022; Huber et al., 2022)RI(Wang et al., 2024a; Olsen et al., 2022; Merrick and Taly, 2020; Redelmeier et al., 2020; Aas et al., 2021b, a; Frye et al., 2021; Kumar et al., 2021b; Covert and Lee, 2020; Mitchell et al., 2022; Jethani et al., 2021; Wang et al., 2023; Luo et al., 2022)(Kumar et al., 2020; Duval and Malliaros, 2021; Kwon and Zou, 2022; Zheng et al., 2022b; Demertzis et al., 2022; Cuzzocrea et al., 2023; Sundararajan and Najmi, 2020; Kolpaczki et al., 2024a; Bento et al., 2021; Ma and Tourani, 2020; Williamson and Feng, 2020; Olsen et al., 2024; Huang and Marques-Silva, 2024)DC(Azcoitia et al., 2022; Tian et al., 2022b; Zhang et al., 2023b; Wu et al., 2023b; Xia et al., 2025)DV(Jia et al., 2019; Zhang et al., 2023a; Shobeiri and Aajami, 2021; Xu et al., 2023; Ghorbani et al., 2020; Kwon et al., 2021; Sun et al., 2024)DV(Schoch et al., 2022; Shobeiri and Aajami, 2022; Xia et al., 2024; Ghorbani and Zou, 2019; Wang et al., 2024d)DM(Courtnage, 2022; Ghorbani et al., 2022; Shim et al., 2021; Courtnage and Smirnov, 2021; Pandl et al., 2021)RI(Heuillet et al., 2022; Ruggeri et al., 2024; Li et al., 2021)RI(Bertossi et al., 2023; Deutch et al., 2022; Livshits et al., 2021b, a)DC(Luo et al., 2024; Xia et al., 2023; Liu et al., 2021; Agarwal et al., 2019; Tian et al., 2022a; Si and Pei, 2024; Sim et al., 2020; Nguyen et al., 2022; Sim et al., 2020)DV(Liu et al., 2023; Sim et al., 2022; Wang et al., 2024b; Garrido Lucero et al., 2024)DM(Zhang et al., 2020)RI(Farchi et al., 2021)RI(Brandsæter and Glad, 2024)RT(Xu et al., 2024)DM(Cheng et al., 2024; Zheng et al., 2022a; Yang et al., 2023b, 2022a; Fan et al., 2022; Yang et al., 2022b, 2024; Liu et al., 2022a, 2020; Tastan et al., 2024; Xu et al., 2021; Wang et al., 2024c)(Arbaoui et al., 2024; Batool et al., 2022; Lin et al., 2023; Yu et al., 2020a; Ma et al., 2021; Huang et al., 2020; Shi et al., 2022; Yu et al., 2020b; Wei et al., 2020; Liu et al., 2022b; Wang et al., 2022a; Zhu et al., 2023b; You et al., 2021; Wang et al., 2022b)DM(Guan et al., 2022; Wang et al., 2022b; Huang et al., 2021; Otmani et al., 2024; Xi et al., 2021; Singhal et al., 2024; Wang et al., 2020; Nagalapatti and Narayanam, 2021; Salarbashishahri et al., 2022; Rozemberczki and Sarkar, 2021)DM(Wu et al., 2023a; Salarbashishahri et al., 2022; Sun et al., 2023; Tang et al., 2021; Yang et al., 2023a; Franses et al., 2024; Drungilas et al., 2023)DM(Bimonte et al., 2024)❶❷❶❷❶❷❶❷❶❷❶❷❶❷
@@ 3.1. Cooperative Game Modeling
Finding 1. Players may differ from the analytical object. While players in SV typically align with the analytical objects of a DA task, they can also represent sub-components of the analytical object for specialized objectives.
@@
In contrast, Wang et al. (2019) defined the players as the sub-components of the analytical object in a multi-group data valuation task. The task is to evaluate contributions to Cervical cancer prediction across five disjoint groups, where each group’s dataset contains three unique data features (e.g., age) with no feature overlap between groups. Here, the player is the data feature, a sub-component of the core analytical object (each group’s dataset). The flexibility in defining players as either the analytical object or its sub-components underscores the adaptability to DA task-specific interpretability needs.
@@
The utility function has two kinds of outputs: (1) the goodness-of-fit score between the DA task’s outputs and ground truth facts, including but not limited to test accuracy, training loss, confidence score, etc., (2) the task outputs themselves, more specifically, predictive outputs from the ML models learned in DA tasks or the answers to queries/questions on the data used in the DA workflow.
@@
Finding 2. Utility definitions are task-dependent. The choice of utility is dictated by the objectives of the DA task. When maximizing or minimizing goodness-of-fit scores (e.g., accuracy), the utility directly reflects those scores. Conversely, when explaining task outputs or quantifying data impacts on those outputs, the utility is defined by the outputs themselves.
@@
Discussion. Figure 1 exemplifies this duality. In the DV task, utility is tied to test accuracy to optimize a medical image classifier, while in the RI task, utility derives from predictive outputs to explain how pixel-level variations influence results. Notably, a single player-utility pairing, such as defining players as ML models and utility as test accuracy, can serve multiple DA tasks (e.g., DM and RT in the figure). Similarly, one task may leverage SV for diverse purposes: the DM task, for instance, can use SV both to weight local models and to assign pricing incentives, ensuring higher model quality while aligning contributions with economic rewards. This adaptability reveals SV’s capacity to address heterogeneous analytical goals through context-aware utility design.
@@
Finding 3. The real-world functionality of applying SV lies in three aspects: (1) to construct fair marketplaces for data and related products, incentivizing data sharing; (2) to improve the quality of DA outputs while reducing the economic cost; (3) to transfer DA outputs into actions to solve real-world problems.
@@
Discussion. Figure 1 exemplifies the three functionalities that different DA tasks target to achieve. DC and RT tasks seek to provide helpful results to construct fair marketplaces for data in the medical domain. DV and DM tasks strive to improve the test accuracy of a medical image classification model while reducing the learning cost. The RI task intends to transfer predictions from a medical image classification model into actions assisting diagnosis. As summarized in Figure 2, to serve the first functionality, the purpose of applying SV is typically pricing. For the second functionality, the purposes of applying SV include pricing, selection, and weighting. To fulfill the last functionality, the purpose of applying SV is primarily attribution. Demanders can design SV applications that satisfy their specific requirements according to Figure 2 and Findings 1-3.